A motorcyclle stunt driver zooms off the end of a cliff at a speed of 30 meters per second. If he lands after 0.75 seconds, what is the height of the cliff?
What is the solution to this problem?
1
Expert's answer
2012-03-22T13:01:55-0400
H = gt²/2 = 9.8*0.75²/2 = 2.75625 m.